# Joint Design of Radar Receive Filter and Unimodular ISAC Waveform with Sidelobe Level Control

### v1

@@ Abstract
Integrated sensing and communication (ISAC) has been considered a key feature of next-generation wireless networks. This paper investigates the joint design of the radar receive filter and dual-functional transmit waveform for the multiple-input multiple-output (MIMO) ISAC system. While optimizing the mean square error (MSE) of the radar receive spatial response and maximizing the achievable rate at the communication receiver, besides the constraints of full-power radar receiving filter and unimodular transmit sequence, we control the maximum range sidelobe level, which is often overlooked in existing ISAC waveform design literature, for better radar imaging performance. To solve the formulated optimization problem with convex and nonconvex constraints, we propose an inexact augmented Lagrangian method (ALM) algorithm. For each subproblem in the proposed inexact ALM algorithm, we custom-design a block successive upper-bound minimization (BSUM) scheme with closed-form solutions for all blocks of variable to enhance the computational efficiency. Convergence analysis shows that the proposed algorithm is guaranteed to provide a stationary and feasible solution. Extensive simulations are performed to investigate the impact of different system parameters on communication and radar imaging performance. Comparison with the existing works shows the superiority of the proposed algorithm.
@@ I Introduction
Integrated sensing and communication (ISAC) unifies sensing and communication (S&\&&C) tasks into a single system, improving efficiency and performance by sharing various resources like spectrum and hardware [1, 2]. With its potential to support emerging applications requiring high-quality wireless connections and accurate sensing, such as autonomous driving and smart homes, ISAC is widely regarded as a key enabler for next-generation wireless networks [3]. However, S&\&&C subsystems have distinct performance requirements [4]: sensing favors unimodular and deterministic waveforms, while communication relies on waveforms with high degrees of freedom (DoFs) and randomness for efficient information transmission. As a result, the interference between them is inevitable.
@@
Since wireless systems often need to simultaneously serve multiple users and meet their various sensing and communication needs, multiple-input multiple-output (MIMO) technology becomes particularly important in ISAC systems. Early ISAC research focused on the MIMO radar-communication coexistence by mitigating interference through spectrum-sharing techniques like dynamic spectrum access [5] and null-space projection [6]. Although these approaches enabled S&\&&C coexistence, the systems were designed separately and required side-information exchange, leading to additional cooperation costs [7]. As an advancement from the spectrum-sharing scheme, dual-functional waveform design has recently drawn a lot of interest [8] for its ability to simultaneously sense targets and transmit information using a single device, eliminating the need for side-information exchange in S&\&&C cooperation [7].
There have been many studies focusing on dual-functional waveform design; see [1] and the references therein. Nevertheless, the majority of these works have not considered controlling the sidelobes of the correlation function. In current communication systems, such as 5G NR, signals usually have a certain bandwidth[9]. Using these signals for ISAC inevitably affects the sensing performance due to range sidelobes. Large sidelobes are undesirable in sensing scenarios because they could interfere with or even mask the responses from targets that are far away or with a small radar cross-section (RCS), which will lead to missed detections [10]. Additionally, even with an optimized radar receive beampattern, high range sidelobe levels still result in false detections. Therefore, sidelobe control is crucial in ISAC waveform design to ensure sensing performance. It is well-known that the power amplifier works most efficiently when the input signal is unimodular [11], thus it is necessary to design the unimodular ISAC waveform. In this paper, we focus on the dual-functional waveform design for the monostatic downlink transmission in MIMO-ISAC system, aiming for maximizing the achievable rate and minimizing the beampattern mean square error (MSE), as well as controlling the peak side-lobe level.
@@ I-A Literature Review
Sensing-centric waveform design focuses on embedding the communication information into sensing waveforms [1]. For example, the communication information can be embedded into chirp waveforms [12], spatial beampattern [13], and ambiguity function [14]. While these approaches generally exhibit strong radar sensing capabilities, they often face low communication rates due to the limited number of embedded information bits. In contrast, communication-centric waveform design implements radar sensing using existing communication waveforms, such as orthogonal frequency division multiplexing (OFDM) [15] and orthogonal time frequency space (OTFS) [16]. However, the sensing performance of communication-centric designs is unpredictable due to the inherent randomness of communication signals and potential distortion from high peak-to-average power ratio.
@@
To address the limitations of separate designs and achieve trade-offs between S&\&&C, many works have focused on joint waveform design [1]. This approach constructs waveforms by solving optimization problems under various S&\&&C constraints. More specifically, in [17], the authors jointly maximized a weighted sum rate while minimizing the radar beampattern approximation MSE, constrained by per-antenna power limits, enabling rate-splitting multiple access and interference management in MIMO-ISAC systems. In [18], the authors minimized the Cramér-Rao bound (CRB) for direction-of-arrival (DoA) estimation by designing the beamforming matrix under individual signal-to-interference-plus-noise ratio (SINR) constraints at each communication receiver and the transmit power budget. Additionally, the work [19] proposed to jointly precode communication and radar waveforms to achieve maximum DoFs in waveform design. Recent progress on joint waveform design in MIMO-ISAC systems has been made in [20] and [21], in which [20] optimized a weighted combination of the sum rate and the CRB for target estimation and [21] maximized the system energy efficiency by constraining the transmit power budget, communication SINR, and target estimation CRB.
@@
The works mentioned previously accomplished a balance between S&\&&C to some extent, but they did not address the issue of range sidelobe control. As for now, there have been few studies on sidelobe control for ISAC systems. The work [10] proposed a MIMO-ISAC waveform design framework that realized an integrated side-lobe level (ISL) reduction by minimizing a weighted sum of beampattern MSE, ISL, and MUI. Alternatively, the work [22] focused on maximizing the SINR at the radar output while ensuring communication performance, which also achieves the ISL reduction. However, the communication channels in the two works are assumed to be frequency-flat fading, whereas the ISAC signals capable of distinguishing symbol-level delays correspond to frequency-selective fading communication channels.
@@
In addition to the ISL metric, the peak side-lobe level (PSL) should be more important for sidelobe control. That’s because the PSL at radar receiver dictates the false alarm probability [23]. High PSL can lead to false alarms [24], making PSL control significant for achieving a low false alarm rate (FAR). Despite its importance, the PSL minimization problem has received relatively little attention in both the radar signal [23] and ISAC waveform design. One of the main difficulties in directly minimizing the PSL is that the design metric is not differentiable and the corresponding optimization problem is a minimax problem. To the best of our knowledge, the existing works on PSL control focus on pure radar sensing, and the PSL control in ISAC scenarios remains unexplored. For instance, the work [11] proposed a two-step scheme to control the PSL in MIMO radar. The work [25] approximated PSL suppression by minimizing an $\ell_{p}$ norm metric with large $p$ for single-antenna radar systems.
@@
Taking the above factors into consideration, the goal of this paper is to design a downlink MIMO-ISAC waveform that maximizes the achievable rate at the communication receivers, makes the radar receiver’s beampattern as close as possible to the desired one, and controls the PSL under the constraints of unimodular transmit sequences. However, as discussed in [11], designing waveforms with good correlation properties under the constraint of unimodular sequences is already a complicated task, and it will be more challenging to further require good communication performance and the desired beampattern. To simplify the optimization problem and obtain more DoFs on waveform design, we consider the joint design of the receive filter and transmit sequence in this paper.
@@ I-B Our Contributions
  • Practical problem formulation: Many practical factors are taken into account in the problem formulation, including beampattern MSE, communication MUI, PSL, and unimodular transmit sequence. An optimization problem is formulated to minimize a weighted sum of the radar receive beampattern MSE and the MUI at the communication receiver. To avoid solving a minimax problem by minimizing the PSL directly, we control the PSL by constraining the level of sidelobe at each range bin. The formulated problem is a large-scale optimization problem with convex and nonconvex constraints.
  • Efficient solution: An inexact augmented Lagrangian method (ALM) algorithm is proposed to solve the formulated problem. Specifically, a block successive upper-bound minimization (BSUM) scheme is custom-designed to solve the subproblems in the ALM algorithm, and the updates for all blocks of variable in the BSUM scheme admits closed-form solutions, which makes the proposed algorithm efficient.
@@
It is worth mentioning that a similar inexact-ALM framework has been proposed in [26]. In [26], radar sensing beamforming is achieved through quantized constant-envelope waveform design while satisfying the communication performance constraints. However, it does not address how to suppress the sidelobes of the correlation function. Low range bin sidelobes and beampattern are equally important for good radar imaging. Furthermore, after considering sidelobe suppression, the convergence analysis of the optimization problem formulated in this paper becomes more complex compared to [26], since the constraint terms introduced by the auxiliary variables in [26] are linear, whereas those in this paper are nonlinear.
@@
Notations: We use $x$, $\mathbf{x}$, $\mathbf{X}$, and $\mathcal{X}$ to represent scalar, column vector, matrix, and set, respectively. The notation vec($\cdot$) represents the vectorization of a matrix by stacking its columns. The notations $\|\cdot\|_{1}$, $\|\cdot\|_{2}$, and $\|\cdot\|_{\text{F}}$ denote correspondingly the $\ell_{1}$, $\ell_{2}$, and Frobenius norms of a matrix, respectively. $\mathbb{C}$ and $\mathbb{R}$ denote the sets of complex and real numbers, respectively; ${}\mathcal{R}\{\cdot\}$ and ${}\mathcal{I}\{\cdot\}$ are the real and imaginary parts of a complex number, respectively. The superscript $(\cdot)^{*}$, $(\cdot)^{\text{T}}$ and $(\cdot)^{\text{H}}$ represent the conjugate, the transpose and the conjugate transpose operations, respectively. $\mathbb{I}_{\mathcal{A}}(A)$ denotes an indicator function of $\mathcal{A}$, and it takes 00 if $A\in\mathcal{A}$ and +$\infty$+ ∞, otherwise. $\otimes$ represents the Kronecker product. $\mathbf{A}_{[i:j,:]}$ represents the sub-matrix from the $i$-th row to the $j$-th row of a matrix $\mathbf{A}$, and $A[i,j]$ represents the element in the $i$-th row and $j$-th column of matrix $\mathbf{A}$. $\mathrm{tr}(\mathbf{A})$ means the trace of a matrix $\mathbf{A}$. $\mathbf{A}\succeq\mathbf{B}$ means $\mathbf{A}-\mathbf{B}$ is positive semidefinite. $[n]$ denotes $\{1,2,\dots,n\}$ for a positive integer $n$.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider an ISAC base station (BS) equipped with a uniform linear array (ULA) of $N_{\text{T}}$ antennas. There are $N_{\text{C}}$ single-antenna communication receivers in the downlink transmission. We consider a colocated radar receive station equipped with a $N_{\text{R}}$-antenna ULA. Below we introduce the mathematical frameworks used in both S&\&&C functions and formulate the optimization problem.
@@ II-A Communication Signal Model
Considering the frequency-selective feature of the channel in the current communication system [9], we implement the channel model proposed in [27], in which the channel between $m$-th receiver and $n$-th BS transmitting antenna, denoted as $\mathbf{H}_{m,n}$, is frequency-selective and assumed to be quasi-stationary during one block transmission, and the discrete-time channel gains are defined by a $q$-ray model. The channel $\mathbf{H}_{m,n}$ can be characterized by a lower triangular Toeplitz matrix with the first column being $[h_{0}^{(m,n)},h_{1}^{(m,n)},\dots,h_{L}^{(m,n)},\mathbf{0}_{T}]^{\text{T}}$, where $L=\lceil\frac{\tau_{\max}}{T_{s}}\rceil$, $T_{s}$ is the duration of one time-slot, $\tau_{\max}$ is the maximum delay spread, and $T$ is the length of the block.
Due to the multi-taps feature of the channel, inter-block interference (IBI) occurs between every two consecutive transmissions (blocks). To eliminate IBI, a cyclic-prefix (CP) of length $L$ is added at each transmission and will be removed at the receiver. The noiseless received signal at the $m$-th receiver is given by $\mathbf{y}_{m}=\sum_{n=1}^{N_{\text{T}}}\tilde{\mathbf{H}}_{m,n}\mathbf{x}_{n}$, where $\tilde{\mathbf{H}}_{m,n}=\bm{\Pi}_{\text{cp}}\mathbf{H}_{m,n}\bm{\Gamma}_{\text{cp}}$ is a $T\times T$ matrix, $\bm{\Gamma}_{\text{cp}}=[\mathbf{I}_{\text{cp}},\mathbf{I}_{T}]^{\text{T}}$ and $\bm{\Pi}_{\text{cp}}=[\mathbf{0}_{T\times L},\mathbf{I}_{T}]$ denote the CP-inducing and CP-removing matrices, respectively, and $\mathbf{I}_{\text{cp}}$ contains the last $L$ columns of a $T$-dimensional identity matrix $\mathbf{I}_{T}$. Then, the received signals at all $N_{\text{C}}$ users can be represented as
@@
$\gamma_{i}=\frac{\mathbb{E}_{\mathbf{s}_{i}}\{\|\mathbf{s}_{i}\|_{2}^{2}\}}{\mathbb{E}_{\mathbf{s}_{i}}\{\|\mathbf{H}_{i}\mathbf{x}-\mathbf{s}\|_{2}^{2}\}+\sigma_{n}^{2}},$ (4)

### v3

@@ Abstract
Integrated sensing and communication (ISAC) has been considered a key feature of next-generation wireless networks. This paper investigates the joint design of the radar receive filter and dual-functional transmit waveform for the multiple-input multiple-output (MIMO) ISAC system. While optimizing the mean square error (MSE) of the radar receive spatial response and maximizing the achievable rate at the communication receiver, besides the constraints of full-power radar receiving filter and unimodular transmit sequence, we control the maximum range sidelobe level, which is often overlooked in existing ISAC waveform design literature, for better radar imaging performance. To solve the formulated optimization problem with convex and nonconvex constraints, we propose an inexact augmented Lagrangian method (ALM) algorithm. For each subproblem in the proposed inexact ALM algorithm, we custom-design a block successive upper-bound minimization (BSUM) scheme with closed-form solutions for all blocks of the variable to enhance the computational efficiency. Convergence analysis shows that the proposed algorithm is guaranteed to provide a stationary and feasible solution. Extensive simulations are performed to investigate the impact of different system parameters on communication and radar imaging performance. Comparison with the existing works shows the superiority of the proposed algorithm.
@@ I Introduction
Integrated sensing and communication (ISAC) unify sensing and communication (S&\&C) tasks into a single system, improving efficiency and performance by sharing various resources like spectrum and hardware [1, 2]. With its potential to support emerging applications requiring high-quality wireless connections and accurate sensing, such as autonomous driving and smart homes, ISAC is widely regarded as a key enabler for next-generation wireless networks [3, 4]. However, S&\&C subsystems have distinct performance requirements [5]: sensing favors unimodular and deterministic waveforms, while communication relies on waveforms with high degrees of freedom (DoFs) and randomness for efficient information transmission. Unimodular waveform here means the modulus of the complex baseband signal is unity at all times. As a result, the interference between them is inevitable.
@@
Since wireless systems often need to simultaneously serve multiple users and meet their various sensing and communication needs, multiple-input multiple-output (MIMO) technology [6, 7] becomes particularly important in ISAC systems. Early ISAC research focused on the MIMO radar-communication coexistence by mitigating interference through spectrum-sharing techniques like dynamic spectrum access [8] and null-space projection [9]. Although these approaches enabled S&\&C coexistence, the systems were designed separately and required side-information exchange, leading to additional cooperation costs [10]. As an advancement from the spectrum-sharing scheme, some works [11, 12] focused on designing or implementing new antenna structures for ISAC purpose, such as the fluid antenna-aided ISAC systems and its combination with reconfigurable intelligent surface. Besides the newly appeared antenna structures, dual-functional waveform design has recently drawn a lot of interest [13] for its ability to sense targets and transmit information using a single device simultaneously, eliminating the need for side-information exchange in S&\&C cooperation [10] and the change of antenna structures.
Many studies have focused on the design of dual-functional waveforms for ISAC; see [1] and the references therein. However, most of these works overlook the crucial aspect of controlling the sidelobes of the correlation function. In modern communication systems such as 5G NR, the transmitted signals typically occupy a certain bandwidth [14]. When these wideband signals are reused for sensing in ISAC systems, the resulting range sidelobes can significantly degrade sensing performance. High sidelobe levels are undesirable because they may interfere with or even obscure the reflections from distant targets or those with small radar cross-sections (RCS), potentially leading to missed detections [15]. Moreover, even when the radar receiver beampattern is well-optimized, excessive range sidelobes can still cause false alarms. Therefore, effective sidelobe control is essential in ISAC waveform design to ensure reliable sensing performance.
Additionally, it is well known that power amplifiers operate most efficiently when the input signals are unimodular [12], which further motivates the need for unimodular waveform design in ISAC systems. In this paper, we address the problem of dual-functional waveform design for monostatic downlink transmission in a MIMO-ISAC system. Our objective is to jointly maximize the achievable communication rate and minimize the mean square error (MSE) of the radar beampattern while simultaneously controlling the peak sidelobe level.
@@ I-A Literature Review
Sensing-centric waveform design focuses on embedding the communication information into sensing waveforms [1]. For example, the communication information can be embedded into chirp waveforms [16], spatial beampattern [17], and ambiguity function [18]. While these approaches generally exhibit strong radar sensing capabilities, they often face low communication rates due to the limited number of embedded information bits. In contrast, communication-centric waveform design implements radar sensing using existing communication waveforms, such as orthogonal frequency division multiplexing (OFDM) [19] and orthogonal time frequency space (OTFS) [20]. However, the sensing performance of communication-centric designs is unpredictable due to the inherent randomness of communication signals and potential distortion from a high peak-to-average power ratio.
@@
To address the limitations of separate designs and achieve trade-offs between S&\&C, many works have focused on joint waveform design [1]. This approach constructs waveforms by solving optimization problems under various S&\&C constraints. More specifically, in [21], the authors jointly maximized a weighted sum rate while minimizing the radar beampattern approximation MSE, constrained by per-antenna power limits, enabling rate-splitting multiple access and interference management in MIMO-ISAC systems. In [22], the authors minimized the Cramér-Rao bound (CRB) for direction-of-arrival (DoA) estimation by designing the beamforming matrix under individual signal-to-interference-plus-noise ratio (SINR) constraints at each communication receiver and the transmit power budget. Additionally, the work [23] proposed to jointly precode communication and radar waveforms to achieve maximum DoFs in waveform design. Recent progress on joint waveform design in MIMO-ISAC systems has been made in [24] and [25], in which [24] optimized a weighted combination of the sum rate and the CRB for target estimation and [25] maximized the system energy efficiency by constraining the transmit power budget, communication SINR, and target estimation CRB.
@@
The works mentioned previously accomplished a balance between S&\&C to some extent, but they did not address the issue of range sidelobe control. As for now, there have been few studies on sidelobe control for ISAC systems. The work [15] proposed a MIMO-ISAC waveform design framework that realized an integrated sidelobe level (ISL) reduction by minimizing a weighted sum of beampattern MSE, ISL, and MUI. Alternatively, the work [26] focused on maximizing the SINR at the radar output while ensuring communication performance, which also achieves the ISL reduction. However, the communication channels in the two works are assumed to be frequency-flat fading, whereas the ISAC signals capable of distinguishing symbol-level delays correspond to frequency-selective fading communication channels.
@@
In addition to the ISL metric, the peak sidelobe level (PSL) should be more important for sidelobe control. That’s because the PSL at the radar receiver dictates the false alarm probability [27]. High PSL can lead to false alarms [28], making PSL control significant for achieving a low false alarm rate (FAR). Despite its importance, the PSL minimization problem has received relatively little attention in both the radar signal [27] and ISAC waveform design. One of the main difficulties in directly minimizing the PSL is that the design metric is not differentiable, and the corresponding optimization problem is a minimax problem. To the best of our knowledge, the existing works on PSL control focus on pure radar sensing, and PSL control in ISAC scenarios remains unexplored. For instance, the work [29] proposed a two-step scheme to control the PSL in MIMO radar. The work [30] approximated PSL suppression by minimizing an $\ell_{p}$ norm metric with large $p$ for single-antenna radar systems.
@@
Taking the above factors into consideration, the goal of this paper is to design a downlink MIMO-ISAC waveform that maximizes the achievable rate at the communication receivers, makes the radar receiver’s beampattern as close as possible to the desired one, and controls the PSL under the constraints of unimodular transmit sequences. However, as discussed in [29], designing waveforms with good correlation properties under the constraint of unimodular sequences is already a complicated task, and it will be more challenging further to require good communication performance and the desired beampattern. To simplify the optimization problem and obtain more DoFs on waveform design, we consider the joint design of the receive filter and transmit sequence in this paper.
@@ I-B Our Contributions
  • Unimodular ISAC waveform design with range sidelobe level control: Many factors are taken into account in the problem formulation, including beampattern MSE, communication MUI, PSL, and unimodular transmit sequence. An optimization problem is formulated to minimize a weighted sum of the radar receive beampattern MSE and the MUI at the communication receiver. To avoid solving a minimax problem by minimizing the PSL directly, we control the PSL by constraining the level of sidelobe at each range bin. The formulated problem is a large-scale optimization problem with convex and nonconvex constraints.
  • Efficient solution: An inexact augmented Lagrangian method (ALM) algorithm is proposed to solve the formulated problem, where each ALM nonconvex subproblem is solved approximately to avoid the huge number of iterations for obtaining an exact stationary point. Specifically, a block successive upper-bound minimization (BSUM) scheme is custom-designed to solve the subproblems in the ALM algorithm, and the updates for all blocks of variable in the BSUM scheme admit closed-form solutions, which makes the proposed algorithm efficient.
@@
It is worth mentioning that a similar inexact-ALM framework has been proposed in [31]. In [31], radar sensing beamforming is achieved through quantized constant-envelope waveform design while satisfying the communication performance constraints. However, it does not address how to suppress the sidelobes of the correlation function. Low-range bin sidelobes and beampattern are equally important for good radar imaging. Furthermore, after considering sidelobe suppression, the convergence analysis of the optimization problem formulated in this paper becomes more complex compared to [31] since the constraint terms introduced by the auxiliary variables in [31] are linear, whereas those in this paper are nonlinear.
@@
Notations: We use $x$, $\mathbf{x}$, $\mathbf{X}$, and $\mathcal{X}$ to represent scalar, column vector, matrix, and set, respectively. The notation vec($\cdot$) represents the vectorization of a matrix by stacking its columns. The notations $\|\cdot\|_{1}$, $\|\cdot\|_{2}$, and $\|\cdot\|_{\text{F}}$ denote correspondingly the $\ell_{1}$, $\ell_{2}$, and Frobenius norms of a matrix, respectively. $\mathbb{C}$ and $\mathbb{R}$ denote the sets of complex and real numbers, respectively; $\mathcal{R}\{\cdot\}$ and $\mathcal{I}\{\cdot\}$ are the real and imaginary parts of a complex number, respectively. The superscript $(\cdot)^{*}$, $(\cdot)^{\text{T}}$ and $(\cdot)^{\text{H}}$ represent the conjugate, the transpose and the conjugate transpose operations, respectively. $\mathbb{I}_{\mathcal{A}}(A)$ denotes an indicator function of $\mathcal{A}$, and it takes 0 if $A\in\mathcal{A}$ and $\infty$, otherwise. $\otimes$ represents the Kronecker product. $\mathbf{A}_{[i:j,:]}$ represents the sub-matrix from the $i$-th row to the $j$-th row of a matrix $\mathbf{A}$, and $A[i,j]$ represents the element in the $i$-th row and $j$-th column of matrix $\mathbf{A}$. $\mathrm{tr}(\mathbf{A})$ means the trace of a matrix $\mathbf{A}$. $\mathbf{A}\succeq\mathbf{B}$ means $\mathbf{A}-\mathbf{B}$ is positive semidefinite. $[n]$ denotes $\{1,2,\dots,n\}$ for a positive integer $n$.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider an ISAC base station (BS) equipped with a uniform linear array (ULA) of $N_{\text{T}}$ antennas. $N_{\text{C}}$ single-antenna communication receivers are in the downlink transmission. We consider a colocated radar receive station with a $N_{\text{R}}$-antenna ULA. Below, we introduce the mathematical frameworks used in both S&\&C functions and formulate the optimization problem.
@@ II-A Communication Signal Model
Considering the frequency-selective feature of the channel in the current communication system, [14], we implement the channel model proposed in [32], in which the channel between $m$-th receiver and $n$-th BS transmitting antenna, denoted as $\mathbf{H}_{m,n}$, is frequency-selective and assumed to be quasi-stationary during one block transmission. A $q$-ray model defines the discrete-time channel gains. The channel $\mathbf{H}_{m,n}$ can be characterized by a lower triangular Toeplitz matrix with the first column being $h_{0}^{(m,n)},h_{1}^{(m,n)},\dots,h_{L}^{(m,n)},\mathbf{0}_{T}]^{\text{T}}$, where $L=\lceil\frac{\tau_{\max}}{T_{s}}\rceil$, $T_{s}$ is the duration of one time-slot, $\tau_{\max}$ is the maximum delay spread, and $T$ is the length of the block.
Due to the multi-taps feature of the channel, inter-block interference (IBI) occurs between every two consecutive transmissions (blocks). To eliminate IBI, a cyclic prefix (CP) of length $L$ is added at each transmission and will be removed at the receiver. The noiseless received signal at the $m$-th receiver is given by $\mathbf{y}_{m}=\sum_{n=1}^{N_{\text{T}}}\tilde{\mathbf{H}}_{m,n}\mathbf{x}_{n}$, where $\tilde{\mathbf{H}}_{m,n}=\bm{\Pi}_{\text{cp}}\mathbf{H}_{m,n}\bm{\Gamma}_{\text{cp}}$ is a $T\times T$ matrix, $\bm{\Gamma}_{\text{cp}}=[\mathbf{I}_{\text{cp}},\mathbf{I}_{T}]^{\text{T}}$ and $\bm{\Pi}_{\text{cp}}=[\mathbf{0}_{T\times L},\mathbf{I}_{T}]$ denote the CP-inducing and CP-removing matrices, respectively, and $\mathbf{I}_{\text{cp}}$ contains the last $L$ columns of a $T$-dimensional identity matrix $\mathbf{I}_{T}$. Then, the received signals at all $N_{\text{C}}$ users can be represented as
@@
We assume that the BS has perfect channel state information (CSI), which allows us to clearly highlight the performance gain and design principles of the proposed system under ideal conditions. The waveform design under partial or statistical CSI, as in [33], will be our future work.
@@
$\gamma_{i}=\frac{\mathbb{E}_{\mathbf{s}_{i}}\{\|\mathbf{s}_{i}\|_{2}^{2}\}}{\mathbb{E}_{\mathbf{s}}}\{\|\mathbf{H}_{i}\mathbf{x}-\mathbf{s}\|_{2}^{2}\}+\sigma_{n}^{2},$ (4)